How many gallons of a liquid that is 74 percent alcohol must be combined with 5 gal of another that is 90 percent alcohol to obtain a mixture that is 84 percent alcohol?

Answer by stanbon(57246)

How many gallons of a liquid that is 74 percent alcohol must be combined with 5 gal of another that is 90 percent alcohol to obtain a mixture that is 84 percent alcohol?
----
Equation:
alcohol + alcohol = alcohol
0.74x + 0.90*5 = 84*(x+5)
-----------------
Multiply thru by 100 to get:
74x + 90*5 = 84x + 84*5
---
10x = 6*5
x = 3 gallons (amt. of 74% solution needed)
==============
Cheers,
Stan H.